OpenStudy (anonymous):

solve each system by substitution Please show how you got it and also how you found Y... Thanks :) 4x+2y=10 x-y=13

OpenStudy (radar):

I would choose one of the equations to be substituted. I would pick the simplest or the x-y=13. I then would choose which variable I am going to substitute. I choose x as x=y+13. Substitute for x in the first equation (4x+2y=10) 4(y+13) + 2y = 10 4y +52 +2y = 10 6y=10-52=-42 y=-7 We solved for y now substitute for y in either equation: (I choose the 2nd one) x-(-7)=13 x+7=13 x=13-7=6